Status Report

ISS On-Orbit Status 6 Feb 2002

By SpaceRef Editor
February 6, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted
previously or below.

After Week 8 got underway with some excitement, excellent progress is
being made by the crew to “catch up”.  Besides offering a full
plate of individual systems and research assignments for them, today
is also CDR Onufrienko?s “Den? rozhdeniya”. Happy Birthday,
Yuri Ivanovich!  On this occasion, he had a private family
conference (PFC) via Ku-band/video and S-band/audio today, followed
by a telephone conference (S-band) with the JSC Astronaut Office.

Onufrienko took less than an hour to outfit the TORU (teleoperator
control system) station in the SM with the worksite for a second
operator.  TORU is used as manual backup for proximity and
docking operations of visiting Progress vehicles in the event their
primary KURS rendezvous/docking radar fails.

He also completed the regular daily tasks of SOSH life support system
maintenance and ASU toilet insert replacement, followed by
termination of regeneration of filter cartridge #1 of the BMP harmful
impurities unit, switching it to purification mode and initiating the
24-hr. bake-out process on absorbent bed #2.

FE-1 Carl Walz set up the HRF (human research facility) workstation
for operation and ran a function check on it, followed by concluding
operations.  Pre- and post-checkup commands for the powered
phase came from the PRO (Payload Rack Officer) at the Huntsville POC
(Payload Operations Center), who is linked by computer to all payload
racks aboard the station. PRO monitors and configures the resources
and environment for science experiments, including EXPRESS racks
(multiple-payload racks designed for commercial payloads).  The
HRF was launched on Mission 5A.1 and is located in the Lab, along
with the other racks.

The crew and the two mission control centers, Houston and Moscow,
joined in integrated  troubleshooting of a tripped remote power
controller (RPC Z13BA), which controls one channel of the electric
power current between the US segment (USOS) and Russian segment (RS).
 US power is provided through the voltage-converting ARCUs
(American-to-Russian converter units), and the checkout, after
initial visual inspection, focused on connectors in the FGB, fed from
the PMA-1 (pressurized mating adapter #1).  After pin checks and
resistance measurements between contacts, it was determined that the
feeder circuit from the US side is OK and that, judging from a low
resistance reading, the problem appears to be in one of two current
stabilizers (ST-25) used to charge the storage batteries of the Soyuz
docked to the FGB.

CDR Onufrienko discussed the IMS (inventory management system) audit
of life support consumables with MCC-Moscow.  As reported, a
discrepancy between IMS database values and actual supplies on board
needs to be cleared up, primarily to assist with improving
assumptions of consumption/depletion rates for planning of future
resupply manifests.

FE-2 Dan Bursch collected air and surface samples using the MAS
(microbial air sampler) and SSK (surface sample kit) equipment.

Both he and FE-1 Carl Walz performed their PFE (periodic fitness
evaluation) tests, each assisting the other in turn.  They also
completed the weekly data collection for the interpersonal relations
“Interactions” experiment (NTXN).

The results of the on-board acoustic noise survey performed by the
crew on 1/29 are now available. Data show that permissible overall
sound pressure levels in the SM and DC-1 work areas were exceeded
during daytime by up to 11.0 dBa (decibel absolute). During nighttime
in the crew cabins they were exceeded by up to 19.0 dBa.
 Noisiest spots in the SM are near the Vozdukh system and SKV-2
air conditioner, in the DC-1 near the EVA hatch. There has been
significant deterioration of the acoustic environment in the crew
cabins compared to previous sound level surveys during Increment 2.
 Crew was advised to continue using noise protection aids
available onboard the station.  Noise reduction/abatement
clearly continues to be a critically needed technology..

An updated EVA timeline for 2/20 was uplinked, along with reference
material, pictures and notes.  As part of preparations, MCC-H
suggests the crew to run through an entire EVA dry-run, as a dress
rehearsal, on 2/15 (Friday before the EVA).  It would take about
half a day and include suit donning, prebreathe prep, EVA prep
including operations at pressure reduced to 10.2 psi, and some
post-EVA procedures.  The final destination of the Russian
Strela-2 adapter (outside the FGB on an EFGF grapple fixture or
inside the station) is still being discussed between Houston and

The EarthKAM equipment, installed yesterday at an SM nadir window,
today downlinked its electronic images throughout the day and night
under ground control.

Targets for the Russian Uragan earth imaging program today were the
coastline of Sumatra Island (nadir and groundtrack right), Kalimantan
Island, and Philippines using the Kodak DSC (focal length 400 mm),
the airplane crash site on the slope of Cumbal Volcano using the
Rubinar binocular telescope, and Guatemala, Mexico, using the LIV
video cam.

The U.S. CEO (crew earth observations) programs today focused on
Angolan Biomass Burning (best viewing conditions for
aerosols this pass were to right of track over the water as ISS
approached the Angolan coast from the southwest), Lake Nasser,
Toshka Lakes; Egypt
(of interest: views of the Nile River and
Lake Nasser to left of track, for documenting conditions along the
shoreline of the lake and note seasonal water color changes in the
river and the lake) Mediterranean Dust and Smog
(atmospheric and lighting conditions offered good opportunities to
observe smog plumes over the Eastern Mediterranean. Crew was to try
for limb shots to the right of track), Parana River
(well-illuminated pass for documenting the large reservoirs the
Brazilians have constructed in the highlands east of the Parana
River), and Eastern United States (high pressure was
moving over and off the eastern seaboard of the US this pass. Crew to
try for limb views to the left of track to detect aerosols moving

U.S. and Russian Segment Status (as of 1:55 pm EST

Environmental Control and Life Support (ECLSS):
Elektron O2 generator is On (16 Amp mode). Vozdukh CO2
scrubber is in MANUAL cycle mode #5 (vacuum pump failed).  U.S.
CDRA CO2 scrubber is Off.
SM Working Compartment:  Pressure (mmHg) — 753, temperature
(deg C) — 25.8, ppO2 (mmHg) — 137.1, ppCO2 (mmHg) — 2.5.
SM Transfer Compartment:  Pressure (mmHg) — 743, temperature
(deg C) — 20.8; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
FGB Cabin:  Pressure (mmHg) — 744, temperature (deg C) — 22.3;
ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
Node:  Pressure (mmHg) — 743.25, temperature (deg C) — 22.6
(shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
U.S. Lab:  Pressure (mmHg) — 743.7, temperature (deg C) —
22.4, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
Joint Airlock:  Pressure (mmHg) —  743.4, temperature (deg
C) — 22.4.
(Partial pressures ppO2 and ppCO2 in U.S. segment (USOS) not
available because MCA [major constituent analyzer] is failed and in
Extended Life mode [= a state that preserves mass spectrometer
vacuum but produces no pp data]).

Electrical Power Systems (EPS):
Beta Gimbal Assembly (BGA) 2B in Autotrack mode, BGA 4B in
Directed position (125 degrees).
SM battery #7 is cycling; all other ( 7) SM batteries are in “partial
charge” mode.
FGB battery #1 is cycling; all other (5) batteries are in “partial
charge” mode.
Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby

Thermal Control Systems:
Air conditioner SKV-1 is Off (Freon leak). SKV-2 is On.

Command & Data Handling Systems:
C&C-2 MDM is prime, C&C-1 is back-up, and C&C-3
is in standby.
GNC-1 MDM is prime; GNC-2 is backup.
SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
SM Central Computer (TsVM): 2 redundant lanes (of 3) operational.

Communications Systems:
S-band is operating nominally, and Ku-band is operating
nominally in open loop pointing mode.
Audio subsystem operating nominally (except for ATU #2).
Video subsystem operating nominally (but VTR2 is
MCOR (medium-rate communications outage recorder) operating

SSRMS/Canadarm2 at Port stow position (on Keep Alive power on
both strings).
RWS (robotics workstations) are Off.

ISS Orbit (as of this morning, 7:19 am EST):
Mean altitude — 387.0 km
Apogee — 391.3 km
Perigee — 382.8 km
Period — 92.3 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.0006276
Orbits per 24-hr. day — 15.60
Decay rate — 395 m (mean) in last 24 hours
Revolutions since FGB/Zarya launch (Nov. ?98) — 18364
Solar Beta Angle — +12.1 deg (magnitude decreasing)
Current Flight Attitude — LVLH +XVV ZLV (local vertical/local
horizontal: +X-axis in velocity vector; Z-axis in local vertical),
with TEA (torque equilibrium attitude [pitch: -10 deg, yaw: -7 deg,
roll: 0 deg]).

For more on ISS orbit, see

SpaceRef staff editor.